OpenStudy (anonymous):

cot(10pi/3)+(cot(16pi/3))/(cos(17pi/6))

zepdrix (zepdrix):

So they gave you a bunch of annoying rotated angles. We need to unwind them so we can see what angle they refer to in one rotation along the unit circle. \(\large \dfrac{10\pi}{3}\) See how the angle is larger than \(\large 2\pi\)? We want to subtract \(\large 2\pi\) from this angle. \(\large \dfrac{10\pi}{3}-2\pi \qquad = \qquad \dfrac{10\pi}{3}-\dfrac{6\pi}{3}\qquad = \qquad \dfrac{4\pi}{3}\) Which tells us,\[\large \cot\left(\frac{10\pi}{3}\right) \qquad = \qquad \cot\left(\frac{4\pi}{3}\right)\]

OpenStudy (anonymous):

btw theyre asking for the answer in improper form

zepdrix (zepdrix):

Hmm I'm not sure what that means. Do they just mean, don't round your answer to a decimal?

OpenStudy (anonymous):

im guessing just get the answer without further simplifying

zepdrix (zepdrix):

Nah I don't think that's what they mean :) You know that 4pi/3 is a special angle. They won't expect you to remember 10pi/3. You have to do a little work to refer it back to your special angle.

zepdrix (zepdrix):

\[\large \cot\left(\frac{4\pi}{3}\right) \qquad = \qquad \frac{1}{\sqrt3}\] Right? I mean, if you want to try and do this without simplifying you can, but that seems rather silly. :D

OpenStudy (anonymous):

ok confused at your answer of it equalling 1/sqrt3

zepdrix (zepdrix):

If you don't remember cotangent, that's ok. We can rewrite it in terms of sines and cosines. \[\large \cot x=\frac{\cos x}{\sin x} \qquad \rightarrow \qquad \frac{\cos\left(\dfrac{4\pi}{3}\right)}{\sin\left(\dfrac{4\pi}{3}\right)}\]Do you remember these special angles?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hello?

zepdrix (zepdrix):

Yes you do recall what those special angles give you? So are you able to simplify the fraction? :o